# Robust Bayesian Model Averaging for Linear Regression Models With Heavy-Tailed Errors

### v1

@@ 6 Conclusion
We conclude this article with some promising and important directions for future work. In this article, we have considered two well-known symmetric members of the family of generalized hyperbolic distributions (Gneiting,, 1997), namely, the hyperbolic and the Student-$t$ distributions. Potentially, the parameter $p$ in the $\mathrm{GIG}(p,a,b)$ distribution, the mixing density for scale mixture of normals, can be given a flexible prior to accommodate a richer family of heavy-tailed error distributions. Another direction is to consider asymmetric error distributions by putting a prior on $\mu$ in (3). Other viable directions include scenarios, where the number of predictors and/or the sample size is much larger than the ones considered here. In a very high-dimensional framework, standard MCMC algorithms can become trapped in local modes. There are many possibilities that include 1) exploring alternative MCMC algorithms that are customized for highly multimodal distributions, 2) simplifying the problem by modifying the goal to estimate the posterior mode(s), instead of the entire posterior distribution, or 3) considering alternative continuous shrinkage priors for variable selection such as the horseshoe prior, to name a few. While a very large sample size is desirable for estimation, it can reduce the speed of algorithms substantially, which can create a different kind of computational challenge. In such scenarios, alternative algorithms based on divide-and-conquer methods or those based on subsampling could prove to be useful. Last but not the least, exploring the stability of the algorithms and results under varying hyperparameter choices is also worth considering.

### v2

@@ 6 Conclusion
We conclude this article with some promising and important directions for future work. In this article, we have considered two well-known symmetric members of the family of generalized hyperbolic distributions (Gneiting, (1997)), namely, the hyperbolic and the Student-$t$ distributions. Potentially, the parameter $p$ in the $\mathrm{GIG}(p,a,b)$ distribution, the mixing density for scale mixture of normals, can be given a flexible prior to accommodate a richer family of heavy-tailed error distributions. Another direction is to consider asymmetric error distributions by putting a prior on $\mu$ in (3). Some authors (Gagnon et al., (2020); Hamura et al., (2022)) have theoretically established posterior robustness. Empirical assessment of the robustness of the HTEM in terms of well known yardsticks of robustness, such as the influence function (Kawakami and Hashimoto, (2023)), may be an interesting question to address in the future. Other viable directions include scenarios, where the number of predictors and/or the sample size is much larger than the ones considered here. In a very high-dimensional framework, standard MCMC algorithms can become trapped in local modes. There are many possibilities that include 1) exploring alternative MCMC algorithms that are customized for highly multimodal distributions, 2) simplifying the problem by modifying the goal to estimate the posterior mode(s), instead of the entire posterior distribution, or 3) considering alternative continuous shrinkage priors for variable selection such as the horseshoe prior, to name a few. While a very large sample size is desirable for estimation, it can reduce the speed of algorithms substantially, which can create a different kind of computational challenge. In such scenarios, alternative algorithms based on divide-and-conquer methods or those based on subsampling could prove to be useful. Last but not the least, exploring the stability of the algorithms and results under varying hyperparameter choices is also worth considering.